### v1

@@ 3.3. Well-posedness of the Bochner Laplacian in $\mathbf{H}^{1}_{t}(\Gamma)$
Let $\boldsymbol{\gamma}:[0,L]\rightarrow\Gamma$ be the $C^{1}$ geodesic connecting $\mathbf{x}$ and $\overline{\mathbf{x}}$, i.e. $\boldsymbol{\gamma}(0)=\mathbf{x}$ and $\boldsymbol{\gamma}(L)=\overline{\mathbf{x}}$. In what follows, for every (scalar, vector or tensor-valued) function $\mathbf{f}$ defined on $\boldsymbol{\gamma}$ we write $\mathbf{f}^{*}:=\mathbf{f}\circ\boldsymbol{\gamma}$. Let $\dot{\boldsymbol{\gamma}}:=\frac{\,{\mathrm{d}}\boldsymbol{\gamma}}{\,{\mathrm{d}}t}$ be the tangent vector to $\boldsymbol{\gamma}$. By the chain rule $\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}=(\nabla_{M}\mathbf{v})^{*}\dot{\boldsymbol{\gamma}}$, whence $\mathbf{P}^{*}\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}=(\nabla_{\Gamma}\mathbf{v})^{*}\dot{\boldsymbol{\gamma}}=\mathbf{0}$ (cf. (2.1)). Consequently,
$\frac{1}{2}\frac{\,{\mathrm{d}}}{\,{\mathrm{d}}t}|\mathbf{v}^{*}|^{2}=\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}\cdot\mathbf{v}^{*}=\left(\mathbf{P}^{*}\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}\right)\cdot\mathbf{v}^{*}=0.$
@@
Lemma 3.8 (Poincaré inequality).
Let $1<p<\infty$ and assume $\Gamma$ is of class $C^{3}$. Then, there is a constant $C>0$ depending only on $\Gamma$ and $p$ such that
@@
This is a consequence of the Peetre–Tartar’s Lemma [27, Lemma A.38] applied to operators $A:=\nabla_{\Gamma}\in\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),\mathbb{L}^{p}(\Gamma))$ and $T:=\textrm{id}:\mathbf{W}^{1,p}_{t}(\Gamma)\rightarrow\mathbf{L}^{p}(\Gamma)$—which are respectively injective and compact by 3.6 and 2.10—and inequality (3.24). ∎
A straightforward application of the Lax–Milgram theorem [30, Theorem 1.1] yields the well-posedness of (1.3).
@@
Assume $\Gamma$ is of class $C^{3}$. For each $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$, there exists a unique solution $\mathbf{u}\in\mathbf{H}^{1}_{t}(\Gamma)$ of (1.3). Moreover, there is a constant $C>0$ depending only on $\Gamma$, such that
@@
3.4. $\mathrm{L}^{p}$-based Sobolev regularity of the Bochner Laplacian
@@
With the well-posedness of (1.3) at hand, we are now ready to establish higher regularity results. The main goal is to establish $\mathbf{W}^{m+2,p}$-regularity for any $p\in(1,\infty)$, as before. In striking contrast to Section 3.2, the arguments presented in this section do not require going through “localization”. Instead, we exploit the nontrivial relation between the Bochner–Laplace and componentwise Laplace–Beltrami operators in order to keep all arguments “intrinsic”, that is, on $\Gamma$.
Lemma 3.10 ($\mathbf{H}^{m+2}$-regularity of the Bochner Laplacian on $\Gamma$).
@@
Let $m\geq 0$ be an integer and assume $\Gamma$ is of class $C^{m+3}$ and $\mathbf{f}\in\mathbf{H}^{m}_{t}(\Gamma)$. Then, the solution $\mathbf{u}\in\mathbf{H}^{1}_{t}(\Gamma)$ of (1.3) belongs to the space $\mathbf{H}^{m+2}(\Gamma)$and there exists a constant $C>0$ depending only on $\Gamma$ and $m$ such that
$\|\mathbf{u}\|_{m+2,\Gamma}\leq C\|\mathbf{f}\|_{m,\Gamma}.$ (3.35)
Moreover, $u$ strongly satisfies (VL) in the sense that
@@
Combining (1.3) with C.3 and noticing that $\mathbf{P}\mathbf{f}=\mathbf{f}$, we deduce that $\mathbf{u}$ satisfies
$\sum_{i=1}^{d+1}\int_{\Gamma}\nabla_{\Gamma}u_{i}\cdot\nabla_{\Gamma}v_{i}=\int_{\Gamma}\mathbf{h}[\mathbf{u}]\cdot\mathbf{v},\qquad\forall\mathbf{v}\in\mathbf{H}^{1}(\Gamma);$
where $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\mathbf{B}^{2}\mathbf{u}+(2\mathbf{B}:\nabla_{M}\mathbf{u}+\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}\in\mathbf{L}^{2}(\Gamma)$ satisfies
[…]
In particular, choosing $\mathbf{v}=\alpha\mathbf{e}_{k}$ for fixed $k=1,\dotsc,d+1$ and arbitrary $\alpha\in\mathrm{H}^{1}(\Gamma)$ yields
$\int_{\Gamma}\nabla_{\Gamma}u_{k}\cdot\nabla_{\Gamma}\alpha=\int_{\Gamma}h_{k}[\mathbf{u}]\alpha,\qquad\forall\alpha\in\mathrm{H}^{1}(\Gamma).$ (3.37)
That is, $u_{k}-\frac{1}{|\Gamma|}\int_{\Gamma}u_{k}=:\overline{u}_{k}\in\mathrm{H}^{1}_{\#}(\Gamma)$ satisfies the weak formulation (1.2) of the Laplace–Beltrami operator with right-hand side $h_{k}[\mathbf{u}]$. Since $\mathbf{B}$ is of class $C^{m+1}$ (because $\Gamma$ is of class $C^{m+3}$), we successively apply 3.1 to problem (3.37) to deduce that $\mathbf{u}\in\mathbf{H}^{m+2}(\Gamma)$ and that estimate (3.35) holds.
Since $\mathbf{u}\in\mathbf{H}^{2}_{t}(\Gamma)$, we can integrate (1.3) by parts using [11, Ap. B] (valid by density) to deduce (3.36). ∎
We next establish the $\mathbf{W}^{m+2,p}$-regularity for $p>2$. The argument relies on the analogous result of 3.2 for the Laplace–Beltrami operator and bootstrapping.
Lemma 3.11 ($\mathbf{W}^{m+2,p}$-regularity ($p>2$) of the Bochner Laplacian on $\Gamma$).
Let $2<p<\infty$ and assume $\Gamma$ is of class $C^{m+3}$ for some nonnegative integer $m$. If $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, then the solution $\mathbf{u}\in\mathbf{H}^{m+2}_{t}(\Gamma)$ of (1.3) and (3.36) (cf. 3.10) belongs to the space $\mathbf{W}^{m+2,p}_{t}(\Gamma)$ and there exists a constant $C>0$ depending only on $\Gamma$, $m$ and $p$ such that
$\|\mathbf{u}\|_{m+2,p;\Gamma}\leq C\|\mathbf{f}\|_{m,p,\Gamma}.$ (3.38)
@@
We proceed via a bootstrapping argument similar to one in the proof of 3.2. We first observe that $\mathbf{B}$ is of class $C^{m+1}$ (because $\Gamma$ is of class $C^{m+3}$). Notice the following Sobolev embeddings: if $\mathbf{u}\in\mathbf{W}^{m+2,\frac{2d}{d-2n}}(\Gamma)$ for some integer $0\leq n<d/2$, then $\mathbf{W}^{m+1,\frac{2d}{d-2n}}(\Gamma)\hookrightarrow\mathbf{W}^{m,q}(\Gamma)$ for all $q\in(1,\infty)$, if $2(n+1)\geq d$ (cf. 2.7); and $\mathbf{W}^{m+1,\frac{2d}{d-2n}}(\Gamma)\hookrightarrow\mathbf{W}^{m,\frac{2d}{d-2(n+1)}}(\Gamma)$, if $2(n+1)\geq d$ (cf. 2.6). These imply that $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\mathbf{B}^{2}\mathbf{u}+(2\mathbf{B}:\nabla_{M}\mathbf{u}+\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}$ from the proof of 3.10 belongs to $\mathbf{W}^{m,q}(\Gamma)$, where
$q:=\begin{cases}p,&\text{if $2(n+1)\geq d$},\\ \min\{p,\frac{2d}{d-2(n+1)}\},&\text{if $2(n+1)<d$},\end{cases}$
with the estimate
$\|\mathbf{h}\|_{m,q;\Gamma}\leq C\left(\|\mathbf{u}\|_{m+2,\frac{2d}{d-2n},\Gamma}+\|\mathbf{f}\|_{m,p,\Gamma}\right).$ (3.39)
Hence, a direct application of 3.2 ($\mathrm{W}^{m+2,p}$-regularity ($p>2$) of the Laplace–Beltrami operator on $\Gamma$) to problem (3.37) yields $\overline{u}_{k}\in\mathrm{W}^{m+2,q}_{\#}(\Gamma)$ for each $k=1,\dotsc,d+1$ and
$\|\overline{u}_{k}\|_{m+2,q;\Gamma}\lesssim\|h_{k}[\mathbf{u}]\|_{m,q;\Gamma}\lesssim\left(\|\mathbf{u}\|_{m+2,\frac{2d}{d-2n},\Gamma}+\|\mathbf{f}\|_{m,p,\Gamma}\right).$
Consequently, $\mathbf{u}\in\mathbf{W}^{m+2,q}(\Gamma)$ with the same bound.
Recalling that $\mathbf{u}\in\mathbf{W}^{m+2,2}(\Gamma)=\mathbf{H}^{m+2}(\Gamma)$ satisfies (3.35), we eventually obtain $q=p$ by iterating over $n$ starting from $n=0$. ∎
For $1<p<2$ we proceed by a “duality” argument as we did for the Laplace–Beltrami operator (cf. (3.11)). Let $p^{*}:=\frac{p}{p-1}\in(2,\infty)$ be the dual exponent of $p$. Given $\mathbf{f}\in(\mathbf{W}^{2,p^{*}}_{t}(\Gamma))^{\prime}$ consider the problem: find $\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ such that
$\int_{\Gamma}\mathbf{u}\cdot(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v})=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma).$ (3.40)
Since the Bochner Laplacian operator $\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is a continuous bijection from $\mathbf{W}^{2,p^{*}}_{t}(\Gamma)$ onto $\mathbf{L}^{p^{*}}_{t}(\Gamma)\equiv(\mathbf{L}^{p}_{t}(\Gamma))^{\prime}$ (cf. 3.11), it follows that $(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma})^{\mathrm{t}}:\mathbf{L}^{p}_{t}(\Gamma)\rightarrow(\mathbf{W}^{2,p^{*}}_{t}(\Gamma))^{\prime}$ defined by $\langle-(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma})^{\mathrm{t}}\mathbf{u},\mathbf{v}\rangle=-\int_{\Gamma}\mathbf{u}\cdot(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v})$ is also a continuous bijection, according to A.1 Therefore, there exists a unique $\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ solution of (3.40); moreover, there exists a constant $C>0$ depending only on $\Gamma$ and $p$ such that
$\|\mathbf{u}\|_{0,p;\Gamma}\leq C\|\mathbf{f}\|_{(\mathbf{W}^{2,p^{*}}_{t}(\Gamma))^{\prime}}.$ (3.41)
For $1<p<2$, the intrinsic approach employed in 3.11 requires having not only the $\mathrm{W}^{m+2,p}$-regularity of the Laplace–Beltrami problem, but also its well-posedness in $\mathrm{W}_{\#}^{1,p}(\Gamma)$ (see the proof of 3.15), which we do not yet have. For that reason, before establishing the $\mathbf{W}^{m+2,p}$-regularity ($1<p<2$) of the Bochner Laplacian, we revisit the Laplace–Beltrami operator and establish the well-posedness of (LB) for $f\in(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$. More precisely, given $f\in(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, we seek $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ such that
The proof reduces to verifying that problem (3.42) satisfies the conditions of A.2 (Banach–Nečas–Babuška theorem); see also [27, Theorem 2.6]. To that end, we first prove the following two auxiliary results.
Proposition 3.12 (injectivity of $\Delta_{\Gamma}$ in $\mathrm{W}^{1,p}_{\#}(\Gamma)$).
Let $1<p<\infty$ and let $p^{*}:=\frac{p}{p-1}$ be the dual exponent of $p$. Assume $\Gamma$ is of class $C^{2}$. If $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ satisfies
$\int_{\Gamma}\nabla_{\Gamma}u\cdot\nabla_{\Gamma}v=0,\qquad\forall v\in\mathrm{W}^{1,p^{*}}_{\#}(\Gamma),$
then $u=0$ a.e. on $\Gamma$.

### v2

@@ 3.1. $\mathrm{L}^{p}$-based Sobolev regularity
The rest of the proof follows almost verbatim to [34, Lemma 2.1], which we describe below for the sake of completeness. Let $\boldsymbol{\gamma}:[0,L]\rightarrow\Gamma$ be the $C^{1}$ geodesic connecting $\mathbf{x}$ and $\overline{\mathbf{x}}$, i.e. $\boldsymbol{\gamma}(0)=\mathbf{x}$ and $\boldsymbol{\gamma}(L)=\overline{\mathbf{x}}$. In what follows, for every (scalar, vector or tensor-valued) function $\mathbf{f}$ defined on $\boldsymbol{\gamma}$ we write $\mathbf{f}^{*}:=\mathbf{f}\circ\boldsymbol{\gamma}$. Let $\dot{\boldsymbol{\gamma}}:=\frac{\,{\mathrm{d}}\boldsymbol{\gamma}}{\,{\mathrm{d}}t}$ be the tangent vector to $\boldsymbol{\gamma}$. By the chain rule $\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}=(\nabla_{M}\mathbf{v})^{*}\dot{\boldsymbol{\gamma}}$, whence $\mathbf{P}^{*}\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}=(\nabla_{\Gamma}\mathbf{v})^{*}\dot{\boldsymbol{\gamma}}=\mathbf{0}$, because $(\nabla_{\Gamma}\mathbf{v})^{*}=\mathbf{0}$. Consequently, the tangent vector $\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}=\mathbf{P}\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}$ vanishes and
$\frac{1}{2}\frac{\,{\mathrm{d}}}{\,{\mathrm{d}}t}|\mathbf{v}^{*}|^{2}=\frac{\,{\mathrm{d}}\mathbf{v}^{*}}{\,{\mathrm{d}}t}\cdot\mathbf{v}^{*}=0.$
@@
In the proof above, we needed $\Gamma$ to be of class $C^{2}$. This requirement shall propagate to all well-posedness and regularity results in this paper, even in the $\mathrm{L}^{2}$-based setting. It would be interesting to determine if such requirement can be relaxed to $\Gamma$ of class $C^{1,1}$. We leave this question open for future research.
@@
Theorem 3.3 (Poincaré inequality).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Then, there is a constant $C>0$ depending only on $\Gamma$ and $p$ such that
@@
This is a consequence of the Peetre–Tartar’s Lemma [26, Lemma A.38] applied to operators $A:=\nabla_{\Gamma}\in\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),\mathbb{L}^{p}(\Gamma))$ and $T:=\textrm{id}:\mathbf{W}^{1,p}_{t}(\Gamma)\rightarrow\mathbf{L}^{p}(\Gamma)$—which are respectively injective and compact by 3.2 and 2.12—and inequality (3.3). ∎
A direct application of the Lax–Milgram theorem [29, Theorem 1.1] yields the well-posedness of (1.2) for $p=2$.
@@
Assume $\Gamma$ is of class $C^{2}$. For each $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$, there exists a unique solution $\mathbf{u}\in\mathbf{H}^{1}_{t}(\Gamma)$ of (3.2) for $p=2$. Moreover, there is a constant $C>0$, depending only on $\Gamma$, such that
@@
With the well-posedness of (1.2) for $p=2$ at hand, we now proceed to establish its well-posedness for any $p\in(1,\infty)$. In striking contrast to the development of the scalar elliptic theory in [5], the arguments presented in this section do not require going through “localization”. Instead, we exploit the nontrivial relation between the Bochner–Laplace and componentwise Laplace–Beltrami operators induced by the identities in B.1 in order to keep all arguments “intrinsic”, that is, on $\Gamma$.
Before establishing the well-posedness of problem (3.2), we point out the following vector-valued analogue of 2.14 (surjectivity of the divergence operator), which is an immediate consequence of i) of [5, Theorem A.1] (characterization of injectivity and surjectivity), the duality identification $(\mathbb{L}^{p}_{t}(\Gamma))^{\prime}\equiv\mathbb{L}^{p^{*}}_{t}(\Gamma)$, and 3.3 (Poincaré inequality).
Proposition 3.5 (surjectivity of the divergence operator).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Let
$\mathbb{L}^{p}_{t}(\Gamma):=\{\mathbf{A}\in\mathbb{L}^{p}(\Gamma):\quad\mathbf{P}\mathbf{A}\mathbf{P}=\mathbf{A},\quad\text{on $\Gamma$}\},$
which is a closed subspace of $\mathbb{L}^{p}(\Gamma)$. Then, the linear and continuous operator $\operatorname{div}_{\Gamma}:\mathbb{L}^{p}_{t}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ given by
$\langle\operatorname{div}_{\Gamma}\mathbf{A},\mathbf{v}\rangle=-\int_{\Gamma}\mathbf{A}:\nabla_{\Gamma}\mathbf{v},\qquad\forall\mathbf{A}\in\mathbb{L}^{p}_{t}(\Gamma),\ \forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
is surjective.
Theorem 3.6 (well-posedness of the Bochner Laplacian in $\mathbf{W}^{1,p}_{t}(\Gamma)$).
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a unique solution $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ of (3.2). Moreover, there exists a constant $C>0$, depending only on $\Gamma$ and $p$, such that
$\|\mathbf{u}\|_{1,p;\Gamma}\leq C\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}.$ (3.9)
Proof.
We first argue for $p\in(2,\infty)$. Since $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\subseteq(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$, we know from 3.4 that there exists a unique $\mathbf{u}\in\mathbf{H}^{1}_{t}(\Gamma)$ such that
Because of $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$, we know from 3.5 (surjectivity of the divergence operator) and [5, Remark A.3] that exists $\mathbf{A}_{\mathbf{f}}\in\mathbb{L}^{p}_{t}(\Gamma)$ such that $\langle\mathbf{f},\mathbf{v}\rangle=-\int_{\Gamma}\mathbf{A}_{\mathbf{f}}:\nabla_{\Gamma}\mathbf{v}$, for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$, and
$\|\mathbf{A}_{\mathbf{f}}\|_{0,p;\Gamma}\lesssim\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}.$ (3.11)
Consequently, we can equivalently state (3.10) as
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}(\mathbf{P}\mathbf{v})=-\int_{\Gamma}\mathbf{A}_{\mathbf{f}}:\nabla_{\Gamma}(\mathbf{P}\mathbf{v}),\quad\forall\,\mathbf{v}\in\mathbf{H}^{1}(\Gamma),$
or, by means of B.1 and a density argument, also as
$\int_{\Gamma}\nabla_{M}\mathbf{u}:\nabla_{M}\mathbf{v}=\int_{\Gamma}\mathbf{h}\cdot\mathbf{v}+\int_{\Gamma}\mathbf{H}:\nabla_{M}\mathbf{v},\qquad\forall\mathbf{v}\in\mathbf{H}^{1}(\Gamma);$
where $\mathbf{h}:=((\mathbf{A}_{\mathbf{f}}+\nabla_{M}\mathbf{u}):\mathbf{B})\boldsymbol{\nu}$ and $\mathbf{H}:=-(\mathbf{A}_{\mathbf{f}}+\boldsymbol{\nu}\mathbf{u}^{T}\mathbf{B})$. In particular, choosing $\mathbf{v}=\alpha\mathbf{e}_{k}$ for fixed $k=1,\dotsc,d+1$ an arbitrary $\alpha\in\mathrm{H}^{1}(\Gamma)$ yields
$\int_{\Gamma}\nabla_{\Gamma}u_{k}:\nabla_{\Gamma}\alpha=\int_{\Gamma}h_{k}\,\alpha+\int_{\Gamma}\mathbf{H}_{k}\cdot\nabla_{\Gamma}\alpha,\qquad\forall\alpha\in\mathrm{H}^{1}(\Gamma),$
where $\mathbf{H}_{k}$ denotes the $k$-th row of $\mathbf{H}$. That is, $\overline{u}_{k}:=u_{k}-\frac{1}{|\Gamma|}\int_{\Gamma}u_{k}\in\mathrm{H}^{1}_{\#}(\Gamma)$ satisfies (2.9) with right-hand side datum $f_{k}\in(\mathrm{H}^{1}_{\#}(\Gamma))^{\prime}$ given by $\langle f_{k},\alpha\rangle:=\int_{\Gamma}h_{k}\,\alpha+\int_{\Gamma}\mathbf{H}_{k}\cdot\nabla_{\Gamma}\alpha$. Since $\mathbf{B}$ and $\boldsymbol{\nu}$ are respectively of class $C^{0}$ and $C^{1}$ and $\mathbf{A}_{\mathbf{f}}\in\mathbb{L}^{p}(\Gamma)$, and $\mathbf{h}$ depends affinely on $\nabla_{M}\mathbf{u}$ and $\mathbf{H}$ linearly on $\mathbf{u}$, we can proceed via a bootstrapping argument based upon Sobolev embeddings (cf. Section 2.1) and 2.16 to deduce that for each $k\in\{1,\dotsc,d+1\}$, $u_{k}\in\mathrm{W}^{1,p}(\Gamma)$ and
[…]
and so $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ with the same bound. Furthermore, by a density argument it is easy to see that $\mathbf{u}$ satisfies (3.2). This finishes the proof for $p\in(2,\infty)$.
Akin to Step 3 in the proof of [5, Theorem 1.1], the proof for $p\in(1,2)$ follows by duality, noting that the operator induced by the left-hand side of (3.10) is formally self-adjoint. ∎
We now prove the higher $\mathrm{L}^{p}$-based regularity of solutions to (3.2).
Theorem 3.7 ($\mathbf{W}^{m+2,p}$-regularity of the Bochner Laplacian on $\Gamma$).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$. If $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, then the solution $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ of (3.2) belongs to the space $\mathbf{W}^{m+2,p}_{t}(\Gamma)$ and there exists a constant $C>0$, depending only on $\Gamma$, $m$ and $p$, such that
$\|\mathbf{u}\|_{m+2,p;\Gamma}\leq C\|\mathbf{f}\|_{m,p;\Gamma}.$ (3.12)
Moreover, $\mathbf{u}$ strongly satisfies (VL) in the sense that
@@
We proceed similarly to the proof of 3.6 above but without using 3.5. The combination of B.1, a density argument, and integration-by-parts formula [11, Lemma B.7] shows that $\mathbf{u}$ satisfies $\int_{\Gamma}\nabla_{M}\mathbf{u}:\nabla_{M}\mathbf{v}=\int_{\Gamma}\mathbf{h}\cdot\mathbf{v}$ for all $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$ where $\mathbf{h}:=\mathbf{f}+\operatorname{div}_{\Gamma}(\boldsymbol{\nu}\mathbf{u}^{T}\mathbf{B})+(\nabla_{M}\mathbf{u}:\mathbf{B})\boldsymbol{\nu}\in\mathbf{L}^{p}_{t}(\Gamma)$. By choosing $\mathbf{v}=\alpha\mathbf{e}_{k}$ for fixed $k\in\{1,\dotsc,d+1\}$ and arbitrary $\alpha\in\mathrm{W}^{1,p}(\Gamma)$, we have that $\int_{\Gamma}\nabla_{\Gamma}u_{k}:\nabla_{\Gamma}\alpha=\int_{\Gamma}h_{k}\,\alpha$ for all $\alpha\in\mathrm{W}^{1,p^{*}}(\Gamma)$. Successive applications of the higher regularity result from 2.15 (Laplace–Beltrami) conclude the proof of (3.12). Finally, since $\mathbf{u}\in\mathbf{W}^{2,p}_{t}(\Gamma)$, we can integrate (3.2) by parts using [11, Ap. B] (valid by density) to deduce (3.13). ∎
Akin to 2.17 (ultra-weak solutions of the Laplace–Beltrami operator on $\Gamma$), we finish this section by establishing the existence and uniqueness of “ultra-weak” solutions for the Bochner Laplacian.
Lemma 3.8 (ultra-weak solutions of the Bochner Laplacian on $\Gamma$).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2,1}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{2,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a unique $\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$ such that
$\int_{\Gamma}\mathbf{u}\cdot(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v})=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma).$
Moreover, there exists a constant $C>0$ depending only on $\Gamma$ and $p$ such that
$\|\mathbf{u}\|_{0,p;\Gamma}\leq C\|\mathbf{f}\|_{(\mathbf{W}^{2,p^{*}}_{t}(\Gamma))^{\prime}}.$
@@
Since $\Delta_{B}:=-\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ is a linear continuous bijection from $\mathbf{W}^{2,p^{*}}_{t}(\Gamma)$ onto $\mathbf{L}^{p^{*}}_{t}(\Gamma)\equiv(\mathbf{L}^{p}_{t}(\Gamma))^{\prime}$ (cf. 3.7), the proof follows by duality similarly to 2.17 (cf. [5, Lemma 4.3]). ∎
4. Tangent Stokes equations
In this section we apply the well-posedness and regularity theory developed in Section 3 for the (VL) together with the Laplace–Beltrami theory from [5] to derive a corresponding theory for the tangent Stokes equations (S). We refer the reader to the introduction of Section 3 for the sketch of the arguments used in this section. We are concerned with the well-posedness of a general variational formulation for the tangent Stokes equations (S). Let $p\in(1,\infty)$. For each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$ consider the problem: find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
Problem (4.1) is of the generalized Babuška–Brezzi-type (see A.1), whose well-posedness hinges on a collection of inf-sup conditions, which we prove below. Notice that, for $p=2$, (4.1) reduces to (1.4), for which the classical (and simpler) Babuška–Brezzi theory in Hilbert spaces [29, Theorem 2.1] applies.
We start by establishing the following inf-sup condition that encodes, for arbitrary $p\in(1,\infty)$, the surjectivity of the covariant divergence $\operatorname{div}_{\Gamma}$ as an operator acting from $\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ onto $(\mathrm{L}^{p}_{\#}(\Gamma))^{\prime}\equiv\mathrm{L}^{p^{*}}_{\#}(\Gamma)$; compare with 2.14 (surjectivity of the divergence operator).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{1,1}$. Then, there is a constant $C>0$, depending on $\Gamma$ and $p$, such that